Thursday, October 31, 2013

Night Log 2013-10-30

SA: Petri
SO: Fred
Others: Deon, Keith, David + HRS crew for a while

----------------------------------------------------------------------
Summary
----------------------------------------------------------------------


No observations due to wind, humidity and cloud.
Most of the night spent doing Fabry-Perot calibrations.

Wednesday, October 30, 2013

Night Log 2013-10-29

SA: Petri
SO: Veronica
Others: Luke, David, Fred

----------------------------------------------------------------------
Summary
----------------------------------------------------------------------


A very nice stable and dark night with seeing varying
between 1.5" and 2". Science, including ToOs, done for:
2013-1-RSA_OTH-012
2013-1-UW-009
2013-1-POL-003 (x2)
2013-1-RSA_OTH-023 (x2)
2013-1-RSA_OTH-024
2013-1-DC-003

Tuesday, October 29, 2013

Night Log 2013-10-28

SA: Petri
SO: Veronica
Others: Luke

----------------------------------------------------------------------
Summary
----------------------------------------------------------------------


Clear dark night with variable seeing, pretty productive, got many high priority blocks done as well as couple of hours of throughput tests with burst mirror. Humidity stayed just below closing limit for the last 2 hours. Data for:
2013-1-POL-003 (x4)
2013-1-RSA_DC_OTH-001
2013-1-RSA_OTH-016 (x2)

Monday, October 28, 2013

Night Log 2013-10-27

SA: Alexei
SO: Thea
Others:

----------------------------------------------------------------------
Summary
----------------------------------------------------------------------


=> Unstable weather. Clouds from the evening.
Started to be better close to the morning.

=> Some data were taken for the following programs:
2013-1-UKSC-008 (P1)

Sunday, October 27, 2013

Night Log 2013-10-26

SA: Alexei
SO: Thea
Others:

----------------------------------------------------------------------
Summary
----------------------------------------------------------------------


=> High humidity and heavy clouds

=> No science data were taken

Friday, October 25, 2013

Night Log 2013-10-24

SA: Alexei
SO: Thea
Others:

----------------------------------------------------------------------
Summary
----------------------------------------------------------------------


=> Rain, high humidity and heavy clouds.

=> No science data were taken

Thursday, October 24, 2013

Night Log 2013-10-23

SA: Alexei
SO: Thea
Others:

----------------------------------------------------------------------
Summary
----------------------------------------------------------------------


=> Unstable weather with some clouds and bad seeing.
Were closed by humidity finally.

=> Some data were taken for the following programs:
2013-1-RSA_OTH-020 (P2)
2013-1-RSA_DC_OTH-001 (P3)
2013-1-RSA_OTH-008 (P3)
2013-1-RSA_OTH-023 (P0)
2013-1-RSA_OTH-026 (P1)
2013-1-AMNH-002 (P4)
2013-1-UNC_RSA_RU-001 (P3)

Wednesday, October 23, 2013

Night Log 2013-10-22

SA: Brent
SO: Veronica
Others: Alexei

----------------------------------------------------------------------
Summary
----------------------------------------------------------------------


Dark and relatively clear at the start evolved into grey and scattered clouds that cleared at times with good seeing. Data taken successfully for

2013-1-RSA_POL-001
2013-1-RSA_OTH-024
2013-1-RSA_OTH-012
2013-1-RU-001
2013-1-UW-008
2013-1-UKSC-008
2013-1-RSA_OTH-016

Tuesday, October 22, 2013

Night Log 2013-10-21

SA: Brent
SO: Veronica
Others:

----------------------------------------------------------------------
Summary
----------------------------------------------------------------------


Night lost to high humidity.

Monday, October 21, 2013

Night Log 2013-10-20

SA: Brent
SO: Veronica
Others: Luke

----------------------------------------------------------------------
Summary
----------------------------------------------------------------------


Rain, high humidity and thick clouds. No science data taken tonight.

Sunday, October 20, 2013

Night Log 2013-10-19

SA: Brent
SO: Veronica
Others: Luke

----------------------------------------------------------------------
Summary
----------------------------------------------------------------------


Bright night with clouds and poor seeing at start, clearing after middle of night but had several payload problems preventing routine HRS observations. Data taken for
2013-1-RSA_POL-001
2013-1-RSA_DC_OTH-001
and HRS commissioning.

Saturday, October 19, 2013

Night Log 2013-10-18

SA: Brent
SO: Veronica
Others: Luke, Francois

----------------------------------------------------------------------
Summary
----------------------------------------------------------------------


Bright night with lots of cloud at start of night and poor seeing (3-4 arcsec) making observing challenging.
Data taken for
2013-1-RSA_POL-001
and HRS commissioning, before closing around 1am due to high humidity.

Friday, October 18, 2013

Night Log 2013-10-17

SA: Brent
SO: Veronica
Others: Luke, Francois, Paul R, Thea & David

----------------------------------------------------------------------
Summary
----------------------------------------------------------------------


Various comms errors prevented alignment at start of night , then ridge cloud settled in nice and snug for the night. No science data taken. Did a few stray light tests for RSS.

Thursday, October 17, 2013

Night Log 2013-10-16

SA: Brent
SO: Veronica
Others: Luke, Eben, Paul R, Francois, Jonathan

----------------------------------------------------------------------
Summary
----------------------------------------------------------------------


A bright night with thin cirrus and telescope structure problems preventing observing past 8:20 pm. Data taken for
2013-1-UKSC_HET_OTH-001
2013-1-RSA_POL-001

Wednesday, October 16, 2013

Night Log 2013-10-15

SA: Encarni
SO: Thea
Others: David dB, Francois, Hitesh, Luke, Brent

----------------------------------------------------------------------
Summary
----------------------------------------------------------------------


Cirrussy night with bright Moon, ~30 km/h wind and OK seeing.

Science data for:
2013-1-POL_RSA-001 (P3) x4

HRS commissioning data for the second half of the night.

Tuesday, October 15, 2013

Night Log 2013-10-14

SA: Encarni
SO: Thea
Others: David dB, Francois, Luke

----------------------------------------------------------------------
Summary
----------------------------------------------------------------------


Cirrussy night with relatively decent seeing. Frustrating end of the night with technical problems preventing us from doing HRS commissioning work.

Obtained science data for:
2013-1-RU-001 (P0)
2013-1-POL_RSA-001 (P3) x2
2013-1-UW-008 (P1) x2
2013-1-UNC_RSA_RU-001 (P2)

Monday, October 14, 2013

A Busy Week Goes By

Another busy week up here - starting off with some mods to the FIF guider to prevent the cable wrap from slipping out of place & causing problems.  A thin plate was attached to act as a retainer for the cables & a few other tweaks were made - thanks Eben & Ockert!


& we got to welcome a new electronics tech in the form of David de Bruyn.  Here he is, learning the ropes & getting his tracker legs - Great hat by the way :)


Those of us that're obsessed with fibre safety will sleep a little easier now that the last exposed section on the top of the tracker has been covered with a nice solid plate.  Dankie Johan!


Time for some shameless Fabry-Perot marketing...  On the last night of David B's week as SA, he shot this awesome H-alpha sequence of a supernova remnant that had been discovered with the XMM X-ray satellite.  Really neat to see different parts of the nebulosity light up as one scans along in 10 angstrom wavelength steps.


As if we haven't had enough tracker excitement lately, Wednesday was declared H4 service day.  This is the bottom left hexapod, as seen from the tracker platform - the rest will get their turns in due course.


It's a bit disconcerting to climb up to install the fibres when one of the hexapods is simply Not There! 


Happily the fibres all went back in without too much fuss.  The sky ones are quite a challenge, but the object fibres are easy enough to reach & poking a small mirror in behind the FIF confirmed that they're all seated properly.


With the HRS guider, the FIF & the fibres all re-installed, & the sky looking great, we were good to go!  Except - our arc exposures showed no lines?!  Turns out the shutter controller wasn't doing enough to drive the shutter so it wasn't opening at all - aaargh!!  Toggling it manually didn't help & so this looked like a show-stopper (this is the spare controller, the original's already been sent back to Durham for repairs).  Enter the heroic CfAI team: offering expert advice, after midnight, via Skype & email - including ways to reconfigure the shutter controller & thereby saving the day!


We got the big screen in the control room set up to display all the remote desktops needed to run the HRS software...


From everything being in pieces earlier in the day, we managed to get on-sky & send light down the fibres.  Here's the exposure meter running on a 7th magnitude star, albeit with thick clouds wafting over...


Absorption lines galore in this high resolution blue frame of HD10180!


SALT Astronomer Encarni sneakily snapped this shot of us at 04:50 on Thursday morning...  While we may look fairly shattered, it'd been a long day & we'd overcome a number of significant obstacles before managing to get data in each of the 4 modes.  So in spite of all the software, shutter, payload & telescope issues, it ended up being a Great night! :)


After several weeks of reasonably good behaviour, Luke was eventually rewarded with a trip up the CCAS tower...  


CCAS is undoubtedly one of the niftiest SALT locations - thanks Francois!


After that we decided to fire up the iodine cell on HRS & see how that works...


Heating a small iodine crystal inside an evacuated glass tube up to about 80 C causes the iodine to sublimate & the light passing through the cell ends up with a rich forest of absorption lines superimposed.  This allows for extremely accurate radial velocity measurements - which will help us to hunt down exoplanets in future!


Here's what a red frame with the iodine cell on looks like - a huge number of absorption lines down at the blue end of the red chip.


While waiting - in vain - for the wind to drop on Friday night, we started running HRS tours.  The admission fee is wearing a silly outfit, having your photo taken & being published on the blog ;)


The weather was infuriatingly bad for 3 consecutive nights so we took all the calibration frames we could stand.  These included some rather funky ones (shown here with the intensity inverted) which had HRS's internal Thorium Argon arc lamp going down the one fibre (upper trace) & the SALT calibration system's quartz-tungsten-halogen flat field lamps going down the other (lower trace).


Eventually we were driven to decorating Luke's night lunch & we decided to pay homage to our far-away & much-missed colleagues back in Durham...


Jürgen!


Ray!


Eggy!


& David!


OK - enough of  that, it's midnight & Luke's hungry...


In fact, anyone with such neat, legible hand-writing surely needs to be regarded with extreme suspicion, if not fear as well.  How on earth did he ever get a PhD with writing like this?!

 

That's is it from me for a while - 6 weeks since we started all this & it's time for a break...  Thanks to the Sutherland Black Eagles for an incredible send-off as I drove down past the cliff yesterday afternoon!


& to the pair of Jackal Buzzards for the escourt out the observatory gate.


Good luck to Luke as he holds the fort until the Durham guys start to reappear in about a week!  Also to Paul & David for the blue camera surgery which I hate to be missing - please take Lots of photos...  See you all again on the 10th of November!

Night Log 2013-10-13

SA: Encarni
SO: Thea
Others: Luke, Francois, Hitesh. Pierre for a midnight visit.

----------------------------------------------------------------------
Summary
----------------------------------------------------------------------


Photometric night with low humidity and wind, and Ok seeing.

Science data for:
2013-1-GU-002 (P3)
2013-1-RSA_DC_OTH-001 (P1)
2013-1-POL_RSA-001 (P3)

And also some throughput data for HRS commissioning.

Sunday, October 13, 2013

Night Log 2013-10-12

SA: Encarni
SO: Thea
Others: David dB, Lisa and Luke.

----------------------------------------------------------------------
Summary
----------------------------------------------------------------------


Partly cloudy but humid night. Did not get to open.

Took some more HRS calibration data.

Saturday, October 12, 2013

Night Log 2013-10-11

SA: Encarni
SO: Thea
Others: Francois, Hitesh, Tim, David dB, Lisa, Luke. Chris Koen and Patrick Woudt for an HRS tour - thanks Lisa! :)

----------------------------------------------------------------------
Summary
----------------------------------------------------------------------


Frustrating night with wind speeds way over the safe limits. Could not open the telescope. :(

Took some HRS and some MR calibration data.

Friday, October 11, 2013

Night Log 2013-10-10

SA: Encarni
SO: Thea
Others: Lisa, Luke, Francois, Hitesh, David dB, Tim.

----------------------------------------------------------------------
Summary
----------------------------------------------------------------------


Totally clouded out tonight. CCAS tests until about 22:00

Thursday, October 10, 2013

Night Log 2013-10-09

SA: Encarni
SO: Thea
Others: David dB, Hitesh, Francois, Tim.. and Lisa, Luke ALL NIGHT! Pierre for a quick visit.

----------------------------------------------------------------------
Summary
----------------------------------------------------------------------


Beatiful gray/dark clear night with 1.2" seeing, with some banks of clouds passing through as soon as we started HRS commissioning work during the second half of the night.

Science data for:
2013-1-UW-005 (P1)
2013-1-IUCAA-004 (P3)
2013-1-UNC_RSA_OTH-001 (P2)

And despite some technical issues and some perfectly timed pesky clouds, we also got some data through HRS for commissioning purposes. And Lisa declares HRS third light. Hurrah! :)

Wednesday, October 9, 2013

Night Log 2013-10-08

SA: David
SO: Fred
Others: Encarni, Luke, Lisa

----------------------------------------------------------------------
Summary
----------------------------------------------------------------------


Fantastic productive night where everything ran smoothly. Stable & photometric conditions with spells of subarcsec seeing. One of the nicest nights I've see in some time.

Completed alot of blocks:
2013-1-IUCAA-004
2013-1-RSA_DC_OTH-001
2013-1-UNC_RSA_RU-001
2013-1-POL_RSA-001
2013-1-RSA_OTH-023
2013-1-RSA_RU_OTH-001
2013-1-RSA_OTH-024
2013-1-DC-003

Tuesday, October 8, 2013

Night Log 2013-10-07

SA: David
SO: Fred
Others: Keith, Luke, Lisa

----------------------------------------------------------------------
Summary
----------------------------------------------------------------------


Beautiful night, though partly plagued by technical issues.

Completed blocks for:
2013-1-RSA_OTH-012
2013-1-RSA_DC_OTH-001

Monday, October 7, 2013

Night Log 2013-10-06

SA: David
SO: Fred
Others: UK visiting amateurs

----------------------------------------------------------------------
Summary
----------------------------------------------------------------------


Clear at start, but poor seeing and rising humidity brought an eraly end to the night by 02:00.
Observations for:
2013-1-UKSC_HET-OTH-001
2013-1-RSA_UW-001
2013-1-GU-002
2013-1-RSA_OTH-012
2013-1-RSA_OTH-008

Sunday, October 6, 2013

The Past Week

The combination of Blogger Fatigue (27 posts in as many, busy, days) & too much going on up here means that we're overdue for a catch-up.  Quite a lot's happened since the First Light post about a week ago so here's the highlights package...  First of all - Hamish in a clean suit (+ beak!) for a full tour of the tank :)


Before our 2nd night on-sky, we needed to tweak the position of the exposure meter to optimally align the images from the various modes.  While working on this, Jürgen noticed that the lens that focuses the image onto the photomultiplier was the wrong way around & hence not quite playing the game.


That was a simple fix & then it was a matter of much frustrating fiddling in the dark to improve the alignment of the fibre ends at the double scramber (for the High Stability Mode).  After all that we could use the exposure meter to check that the situation was indeed vastly improved for both the HS mode fibres.


During the course of Sunday night (29 September), David (Buckley) decided that pancakes would hit the spot & provide a fine send-off for David (Bramall) & Jürgen who were leavng the next day (courtesy of the overworked Luke Tyas Airport Shuttle).   Dr Buckley devised this crafty solution to address the SALT kitchen's dire need for an egg-beater.


Perhaps this treatment of an electric drill serves to counteract Eben's previous abuse of a smoothie machine in the process of making of a batch of black powder?  Either way, the pancakes were excellent - thanks David!

video

Plenty of progress has been made with the software & so we're now able to take simultaneous images in both the red & the blue channels using the HRS LabVIEW GUIs.  However, we were dismayed to find that the blue camera has undergone a significant focus change for no apparent reason & so this issue remains under investigation.  We also struggled with poor seeing, FIF guider problems & eventually the payload software crashed on us as well, so Not a great night!


More seriously though - the news the next morning (Monday) was that the right-hand Y bearing on the tracker turns out to be badly damaged & hence the telescope would be out of action until a miracle could be arranged.  Fortunately, the Tech Ops team is well versed in miracle production & pretty soon a dramatic plan had been hatched & was being executed.  This involved amputating the lower section of the damaged rail so that the ruptured bearing could be extracted.


Here's Ray eyeing the situation, noting the relative positions of the fibre cable & the electric hacksaw that was used to cut through the slide.


Mechanicals Strydom & Wiid in action!


Cleaning the edge up with a dremel tool & a vacuum cleaner...  One can't overstate the importance of wearing safety specs at a time like this, right Eben?


& there you have it - that extra piece of rail that was only ever going to be in the way.  With this "upgrade", the right-hand Y bearing is now & forever serviceable!


The damaged bearing (just to the left of the vertical plate) could then be removed.


Here it is, along with the extra piece of rail, on the table in the SALT kitchen.


A closer look reveals where the ball bearings have been escaping from!  Very Very ugly & scary to imagine how these 8 mm balls could've popped out while the tracker was somewhere over the primary mirror...


Thanks to Keith for sharing Tech Ops wisdom about how to go about delicate tasks up on the payload.  Apparently tools such as 2.5 mm allen keys, or verniers as shown here, can be held between one's nose & upper lip while using arms & legs to prop oneself in the appropriate position.


This was valuable information as the HRS guider needs to be removed for some modifications so we would need to detach all the fibres from the FIF before heading for Cape Town on Thursday morning.  Each fibre would receive a label & a protective cap before being stored in the fibre holster next to the FIF.


For HRS II we'll figure out some clever, painless way to attach/detach the fibres!


The sky fibres for the High & Medium Resolution modes pose a particular challenge & were awarded "special" names during the extraction process.


All the fibres now have caps on & are safely tucked away.  That's it for now... what remained of the HRS team (Ray, Luke & I) drove down to Cape Town the next day.


We stopped for a Fantastic lunch at Glen Carlou on the way down before dropping Ray off at the airport.  Travel safely Ray & happy teaching, we'll see you out here again in November!


The tracker was back online by Thursday evening - Hugely Well Done guys!!  That's really Great stuff :)  Luke & I will go back up to Suthers on Monday with the bits to fix the guider & to keep going with various commissioning tasks & tests.

Night Log 2013-10-05

SA: David
SO: Fred
Others: Nobody!

----------------------------------------------------------------------
Summary
----------------------------------------------------------------------


Clear to start, but steadily rising RH & plunging temp led to ridge cloud just before midnight.
Short BVIT obs for 2013-1-RSA_HET_UKSC_OTH-001
Completed 2013-1-AMNH-004
Attempted 2013-1-HET-004, but offsetting issue cut obs short
Started 2013-1-RSA_RU_OTH-001, but dewpoint limiit.
Did F-P tests after closing up.

Friday, October 4, 2013

Night Log 2013-10-03

SA: David Buckley
SO: Fred
Others: Ocket & Chris in the beginning

----------------------------------------------------------------------
Summary
----------------------------------------------------------------------


Pointing tests with repaired tracker shows no ill effects as far as pointing & tracking.

Attempted P0 BVIT target (2013-1-RSA_HET_UKSC-001), but problems with fold mirror meant loss of track time.

F-P LR calibrations done successfully & P2 L-R proposal (2013-1-RSA_OTH-013) attempted with partial success.

Thursday, October 3, 2013

Night Log 2013-10-02

SA: Amanda
SO: Fred
Others: Lisa, Ray, Luke, Ockert, Eben, Keith, Chris

----------------------------------------------------------------------
Summary
----------------------------------------------------------------------


Work continued ion the tracker into evening with the successful removal of damaged bearing, which will be replaced tomorrow. hopeful to be back on sky tomorrow night.

HRS testing with calibration arcs continued and the fibres were removed from the FIF for its removal tomorrow to modiify guidance system.

Wednesday, October 2, 2013

Night Log 2013-10-01

SA: Amanda
SO: Veronica
Others: David, Fred, HRS team

----------------------------------------------------------------------
Summary
----------------------------------------------------------------------


. Telescope offline tonight due to tracker problem.

. Data taken for FP calbrations.

Tuesday, October 1, 2013

Night Log 2013-09-30

SA: Amanda
SO: Veronica
Others: HRS team, Janus, David, Hamish

----------------------------------------------------------------------
Summary
----------------------------------------------------------------------


. Telescope offline due to tracker bearing issue.

. Some FP calibrations taken.